OpenStudy (anonymous):

IF YOU ANSWER SOON I WILL MEDAL AND FAN Sarah's mean number of points for four video games is 7,500. How many points must she score in the 5th game so that her mean becomes 7,700

OpenStudy (anonymous):

explain your answer please

OpenStudy (anonymous):

\[7700=\frac{ 7500*4+x }{ 5 }\] find x

OpenStudy (anonymous):

7500 times 4 is 30,000 divided by 5 is 6,000?

OpenStudy (anonymous):

am I correct?

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

can you guys check my answer

OpenStudy (apsara):

your right

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

I can't medal because I already did but I will fan

OpenStudy (anonymous):

7700*5=7500*4+x x=?

OpenStudy (anonymous):

7700x 5 is 38,500 and 7500 times 4 is 30,000 plus 6,000= 36,000? @surjithayer

OpenStudy (anonymous):

38500=30,000+x x=38500-30000=?

OpenStudy (anonymous):

8,500!

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

:) thanks

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

yw